Use logarithmic differentiation to solve dy over dy, if y = xsquared square root 7x-14 over (1+x) to the power 4
1
Expert's answer
2020-06-11T20:43:34-0400

y = "\frac { x²\sqrt {7x-14}} {(1+x)⁴}"

ln(y) = ln [ "\frac { x²\sqrt {7x-14}} {(1+x)⁴}" ]

Using properties of logarithm

ln(y) = ln(x²) + ln "\sqrt {7x-14}" - ln(1+x)⁴

=> ln(y) = 2 ln x + "\frac {1} {2}" ln (7x-14) - 4ln(1+x)

Differentiating with respect to x

"\frac {1} {y}" "\frac {dy} {dx}" = "\frac {2} {x}" + "\frac {7} {2(7x-14)}" - "\frac {4} {1+x}"

So

"\frac {dy} {dx}" = y [ "\frac {2} {x}" + "\frac {7} {2(7x-14)}" - "\frac {4} {1+x}" ]

= "\frac { x²\sqrt {7x-14}} {(1+x)⁴}" [ "\frac {2} {x}" + "\frac {7} {2(7x-14)}" - "\frac {4} {1+x}" ]